# Grounded Test-Time Adaptation for LLM Agents

### v3

@@ 5 Limitations and Future Work
Our study is primarily focused on the Qwen2.5 model family for parametric adaptation; future work should validate these findings across a broader range of open-source architectures. Our non-parametric adaptation method currently operates in environment with explicit state transitions, where testbeds such as user-oriented conversational task is not explored. Additionally, our online parametric adaptation method does not currently normalize updates by the hidden dimension size, which could improve robustness to hyperparameters.
@@ 6 Conclusions
In this work, we presented a comparative study of two distinct strategies for adapting LLM agents to novel, complex environments. We identified two primary failure modes—syntactic and semantic mismatches—and investigated targeted solutions for each. The first, parametric test-time adaptation, is an efficient technique for aligning an agent’s output distribution to the local format of an environment. The second, non-parametric test-time adaptation, is a effective and efficient method for providing the agent with a causal world model through proactive dynamics extraction.
@@ A.1 Experiment details
We follow the BrowserGym guide [^5] https://github.com/ServiceNow/BrowserGym/blob/main/browsergym/webarena/README.md to setup WebArena enviornment. We use AWS EC2 instance with a t3a.xlarge instance. For all exploration episodes of non-parametric test-time adaptation, we use a temperatue of 1.0. We use a temperature of 0.0 for all task evaluation experiments.
@@
We adopt the codebase form (prabhakar_apigen-mt_2025) and use Best-of-N (N=5) strategy for LLM-simulated user agent, so each trajectory will be scored by an LLM judge and the reponse that best adheres with user intent will be select, therefore drastically mitigating variance in average reward.
@@ A.4 Additional results
Table 6: Success rate (SR) of parametric adaptation across different Qwen2.5 instruct model sizes (7B, 14B, 32B) and training hyperparameters. Results are shown for varying numbers of training iterations and two learning rates (LR=0.1 and LR=1.0). Parametric adaptation generally improves SR over the baseline, with larger models and moderate learning rates yielding higher gains.
@@
Table 7: Number of environment dynamics per website before and after filtering with GPT4.1 as the exploration policy and dynamics extractor. Most environment dynamics are filtered out.
@@ A.5 Environment dynamics analysis
Here we provide some concrete examples of environment dynamics that are removed by our filtering mechanism. Exploration policy and dynamics extractor is GPT4.1.

### v4

@@ 5 Limitations and Future Work
Our study is primarily focused on the Qwen2.5 model family for syntactic alignment; future work should validate these findings across a broader range of open-source architectures. Our dynamics grounding method currently operates in environments with explicit state transitions, where testbeds such as user-oriented conversational tasks are not explored. Additionally, our online syntactic alignment method does not currently normalize updates by the hidden dimension size, which could improve robustness to hyperparameters.
@@ 6 Conclusions
In this work, we presented a comparative study of two distinct strategies for adapting LLM agents to novel, complex environments. We identified two primary failure modes—syntactic and semantic mismatches—and investigated targeted solutions for each. The first, syntactic alignment (SA), is an efficient technique for aligning an agent’s output distribution to the local format of an environment. The second, dynamics grounding (DG), is an effective and efficient method for providing the agent with a causal world model through proactive dynamics extraction.
@@
Acknowledgments
We thank Salesforce AI Research for supporting this work. We also thank the reviewers and the area chairs for their time and thoughtful feedback.
@@ A.1 Experiment details
We follow the BrowserGym guide [^5] https://github.com/ServiceNow/BrowserGym/blob/main/browsergym/webarena/README.md to setup WebArena enviornment. We use AWS EC2 instance with a t3a.xlarge instance. For all exploration episodes of dynamics grounding, we use a temperatue of 1.0. We use a temperature of 0.0 for all task evaluation experiments.
@@
We adopt the codebase form (Prabhakar et al., 2025) and use Best-of-N (N=5) strategy for LLM-simulated user agent, so each trajectory will be scored by an LLM judge and the reponse that best adheres with user intent will be select, therefore drastically mitigating variance in average reward.
@@ A.4 Additional results
Table 6: Success rate (SR) of syntactic alignment across different Qwen2.5 instruct model sizes (7B, 14B, 32B) and training hyperparameters. Results are shown for varying numbers of training iterations and two learning rates (LR=0.1 and LR=1.0). Syntactic alignment adaptation generally improves SR over the baseline, with larger models and moderate learning rates yielding higher gains.
@@
Table 7: Number of environment dynamics per website before and after filtering with GPT-4.1 as the exploration policy and dynamics extractor. Most environment dynamics are filtered out.
@@ A.5 Environment dynamics analysis
Here we provide some concrete examples of environment dynamics that are removed by our filtering mechanism. Exploration policy and dynamics extractor is GPT-4.1.